OpenStudy (anonymous):

choose the equation of the line with slope =-5/2; y intercept is (0,7)

OpenStudy (anonymous):

y=-5/2x=7 or y+5/2x-7 or y+5/2x=7 or y+-5/2x-7

OpenStudy (anonymous):

the general equation of a line is y=mx +c where m is the gradient/slope of a line and c is the y intercept of the line. Since slope =-5/2 and intercept = (0,7); y=-5/2x + 7 which is the same as y+5/2x=7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem